Evening Shift August 19

B. Cox (cox@fnal.gov)
Fri, 20 Aug 1999 00:11:08 -0500 (CDT)

Shift Summary for Evening Shift, Aug. 19,/99

On Shift: Brad, Leo

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 3.25
Hours of E832 data (special runs not included): ~6.4
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

16:00 - 16:30 Trigger problem with flaky kumquat input
22:20 - 24:00 Pipeline crate 3 problems
========================================================================
Runs ended this shift:

QK*288-290

=======================================================================

Tapes taken to FCC this shift

QK*288-290

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Low beam intensity for most of shift.

========================================================================
Description of Shift:

15:55 Intensity a bit low. Leo calls MCR. They are still tuning

16:00 ktevana errors - ran kfrend

16:00 - 16:30 trigger problems Run 14276 spills 160-180 approximately

level 2 processor was sending
out a glitch and hanging up.
Problem tracked to the kq AZ lat by Rick.
Unplugged 1y and 2y and error still present.
Unplugged kumquat 3y and 4y and the
the trigger problem went away. When
they were reinstalled the problem stayed
away.

17:50 Switchyard aborts

17:55 Vertical beam too large. Ask MCR to check.

18:55 Beam off horizonally at NM2seed . Call MCR to request they
look into it.

19:25 MCR calls. They plan to reduce intensity for a while to
try to tune up switchyard in an attempt to raise intensity
to the requested level. They will call back when they
complete the tune. They reduce the intensity to 3.5x10**12.

21:15 MCR calls. They are ready to increase the intensity.

21:20 Beam back at close to request of 6.8x10**12

21:45 SEED2 has a slight horizonal roll. Call main corntrol.
They are still tuning.

22:00 MCR finishes. Beam somewhat better but slight roll still
present in the horizonal.

22:20 Run 14276 ends; tapes QK*288-290 full

22:55 Trying to start new run...
Experiencing pipeline problems crate 3

22:55 - 24:00 Struggle with pipeline ends shortly after midnight
Problem solved by remotely power cycling the crate
and rebooting.

========================================================================
Follow up Notes on Previous Shift